Every year, the Parker brothers get together. The combined weight of the Parker brothers is 822 pounds. Bob weighs 12 pounds more than Andy, Carl weighs 17 pounds more than Bob, Dave weighs 23 pounds more than Carl and Ed weighs 207 pounds. How much does Andy weigh?
Answer by JulietG(1812)

Here's what we know:
B = A + 12
C = B + 17 (which is A + 29)
D = C + 23 (which is A + 52)
E = 207
A + B + C + D + E = 822
.
Let's subtract Ed first.
822 - 207 = 615
The equation is now:
A + B + C + D = 615
Substitute the values from above, using the As.
A + (A+12) + (A+29) + (A+52) = 615
Add
4A + 93 = 615
Subtract 93 from each side
4A = 522
Divide each side by 4
A = 130.5
.
therefore B = 142.5, C = 159.5, D = 182.5
.
Let's add them together to prove it.
130.5 + 142.5 + 159.5 + 182.5 + 207 = 822
Success!